OpenStudy (anonymous):

Find the magnitude and direction angle θ, to the nearest tenth of a degree, for the given vector v. v = -4i - 3j

OpenStudy (anonymous):

\[\tan ^{-1} \frac{ -3 }{ -4} =\Theta\ =36.8698976\] so rounding that it's about 36.9 degrees. It's magnitude is found by the pythagorean theorm... so what is that? a^2 + b^2 = c^2 16 + 9 = c^2 c= +- 5 so magnitude is always positive so the answer would be 5@36.9deg